A charged particle of mass 2mg hangs in air in an electric field of strength 4v/m directed upward. Calculate the charge and determine its nature.

Grade:12

1 Answers

Khimraj
6 years ago
For this condition electric force should be balanced by gravity force so nature of charge should be positive. And mg = qE. So q*4 = 0.002*10 so q = 5mC. Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer